# {adjustwidth}0cm-0.2cm Higgs-Induced Gravitational Waves: the Interplay of Non-Minimal Couplings, Kination and Top Quark Mass

### v1

@@ Abstract
We explore a minimal scenario where the sole Standard-Model Higgs is responsible for reheating the Universe after inflation, produces a significant background of gravitational waves and maintains the full classical stability of the electroweak vacuum. As the Higgs self-coupling runs toward negative values at high energy scales, a non-minimal interaction with curvature during a stiff background expansion era drives the Higgs fluctuations closer to the instability scale. This curvature-induced tachyonic instability leads to an intense production of Higgs particles, accompanied by a stochastic gravitational-wave background. The characteristic features of such signal can be directly correlated to the inflationary scale, the non-minimal coupling parameter and the top quark Yukawa coupling. We distinguish between three possible scenarios: absolute stability with low top quark masses, potential vacuum instability, and absolute stability with new physics above the instability scale. Our findings suggest that the detection of a peaked background of gravitational waves together with its inflationary tail has the potential to unveil the features of the Higgs effective potential at very high energy scales while providing a minimal explanation for the reheating phase and the emergence of the Standard-Model plasma in the early Universe. Unlike other studies in the literature, the generation of gravitational waves in our scenario does not depend on the quantum instability of the Standard Model vacuum.
@@ 1 Introduction
In recent years, the steady advancement of particle-collider experiments at the Large Hadron Collider has explored the Higgs sector up to energies of $13.6\times 10^{3}\;\rm GeV$. The several orders of magnitude separating this energy range from plausible inflationary scales up to $10^{16}\textrm{ GeV}$ leave us wondering about new physical phenomena in the Higgs sector that might have taken place in the early Universe. What role does the Higgs play in the history of the early Universe? One possibility is that it dominates the energy budget at very high scales. This assumption identifies the Higgs field with the scalar degree of freedom responsible for inflation, giving rise to Higgs inflation  Bezrukov:2007ep ; Bauer:2008zj ; Garcia-Bellido:2008ycs ; Barbon:2009ya ; Bezrukov:2009db ; Bauer:2010jg ; Burgess:2010zq ; Bezrukov:2010jz ; Giudice:2010ka ; Bezrukov:2014bra ; Hamada:2014wna ; George:2015nza ; Repond:2016sol ; Fumagalli:2016lls ; Bezrukov:2017dyv ; Rubio:2019ypq ; Shaposhnikov:2020fdv ; Dux:2022kuk ; Poisson:2023tja (for a review see Rubio:2018ogq ). In the present work, we will consider the opposite alternative, namely that of a spectator Higgs within the early Universe Markkanen:2017dlc ; Markkanen:2017edu ; Herranen:2014cua ; Herranen:2015ima ; Figueroa:2015rqa ; Opferkuch:2019zbd ; Laverda:2024qjt . In other words, the Higgs remains an energetically-subdominant quantum field during and shortly after inflation with a negligible backreaction on the gravitational background. In general, the whole Standard Model (SM) sector coupled to the Higgs can be considered to be in its vacuum state, while the Universe is still far from achieving a thermalised state.
@@
In this setting, a non-minimal gravitational interaction can introduce significant new-physics effects in the Higgs sector. Indeed, the study of quantum fields on curved spacetimes leads to the conclusion that tree-level operators describing non-minimal gravitational interaction terms are generated naturally when renormalising the field’s energy-momentum tensor Birrell:1982ix . Setting their coupling constants to vanish is an ad-hoc statement that depends on the energy scale. With these considerations, the physics of the Higgs field in the early Universe is dictated by the effective Lagrangian density
@@
$\mathcal{L}=\left(\frac{M^{2}_{P}}{2}-\xi H^{\dagger}H\right)R-g^{\mu\nu}(D_{\mu}H)^{\dagger}(D_{\nu}H)-\lambda\left(H^{\dagger}H-\frac{v^{2}_{{}_{\textrm{EW}}}}{2}\right)^{2}+\mathcal{L}_{\rm SM}+\mathcal{L}_{\rm BSM}\,,$ (1)
where the effective potential of the Higgs doublet $H$ depend on the quartic self-coupling $\lambda$ and the electroweak vacuum expectation value $v_{{}_{\textrm{EW}}}$. The additional term $\mathcal{L}_{\rm SM}$ contains all the SM fields beside the Higgs-field kinetic and potential terms and $\mathcal{L}_{\rm BSM}$ includes potential inputs from physics Beyond the Standard-Model (BSM) at high energy scales. For a Friedmann–Lemaître–Robertson–Walker (FLRW) metric $g_{\mu\nu}={\rm diag}(-1,\,a^{2}(t)\delta_{ij})$, the gravitational contribution to the Higgs mass is specified by the Ricci (or curvature) scalar $R=3(1-3w)\mathcal{H}^{2}$, with $\mathcal{H}=\dot{a}/a$ the Hubble function. Consequently, a variation of the background expansion rate [^1] In the present work it will not be necessary to embed the Higgs sector in a specific inflationary model. The only requirement is the quick onset of a phase of kination after inflation. We therefore maintain an agnostic and general approach to the inflaton dynamics. of the Universe impacts the Higgs effective mass, making it possible for new phases to appear in the effective potential. This is the case for a period of stiff expansion following inflation, since the curvature scalar induces a large positive contribution for $w=-1$ and a large negative contribution for $w>1/3$. For the remainder of this work, we will consider the case of kinetic domination to follow inflation when the departure from slow-rolling leads to a kinetically-dominated Universe with ${w=+1}$. Such a sequence of expansion phases can be found in a plethora of models including asymmetric $\alpha$-attractors Akrami:2017cir ; Dimopoulos:2017zvq ; Dimopoulos:2017tud ; Garcia-Garcia:2018hlc , variable gravity settings Wetterich:1987fm ; Wetterich:1994bg ; Rubio:2017gty , axion-like dynamics Gouttenoire:2021jhk and non-oscillatory quintessential inflation Peebles:1998qn ; Spokoiny:1993kt ; Bettoni:2021zhq . The sign change of the Higgs effective mass at the transition between the two epoch triggers a so-called Hubble-Induced Phase Transition (HIPT) Laverda:2023uqv ; Laverda:2024qjt ; Bettoni:2019dcw ; Bettoni:2021zhq ; Bettoni:2018utf ; Bettoni:2018pbl ; Mantziris:2024uzz ; Kierkla:2023uzo .
Figure 1: Higgs effective potential with a non-minimal gravitational coupling at the onset of kination for three prototypical scenarios: absolute stability (green line, $m_{t}=171.0\textrm{ GeV}$), instability (red line, $m_{t}=171.3\textrm{ GeV}$), stability with new physics as sextic operators at a typical scale ${\Lambda=10^{15}\textrm{ GeV}}$ (blue line, $m_{t}=171.0\textrm{ GeV}$). The cut-off of the effective gravitational theory is given by $\Lambda_{\xi}=9.2\times 10^{15}\textrm{ GeV}$. The cosmological parameters have been set to $\nu=20$ and $\mathcal{H}_{\rm kin}=10^{9}\textrm{ GeV}$.
@@
During inflation, the large gravitational mass suppresses the Higgs quantum fluctuations, both super-horizon and sub-horizon, while preventing the generation of large isocurvature fluctuations Laverda:2024qjt ; Herranen:2014cua ; Herranen:2015ima ; Kohri:2016wof . At the onset of kination, the large negative gravitational mass term induces a spontaneous symmetry-breaking with Hubble-scale-dependent true vacua. The Higgs quantum modes experience an exponential tachyonic amplification that quickly builds up the occupation number of infrared (IR) modes, followed by a cascade effect that turbulently populates ultraviolet (UV) modes as well.  [^2] It is worth noting that this dynamic is significantly different from the well-known scenario of a resonantly-excited Higgs field interacting with an oscillating background Figueroa:2015rqa ; Figueroa:2016dsc ; Figueroa:2017slm ; Figueroa:2016wxr ; Herranen:2015ima ; Mantziris:2020rzh ; Mantziris:2021oah ; Mantziris:2021zox ; Mantziris:2022bfe ; Mantziris:2022fuu ; Mantziris:2023xsp ; Kohri:2016wof ; Postma:2017hbk ; Ema:2017loe ; Enqvist:2016mqj ; Ema:2016kpf . It is precisely these amplified fluctuations that provide us with a way to study the post-inflationary spectator Higgs as they lead to three main outcomes:
@@
    The Universe can be heated by the non-perturbative tachyonic particle production and the corresponding exponential increase in the Higgs energy density. Requiring the Universe to be radiation-dominated at temperatures higher than the Big-Bang-Nucleosynthesis (BBN) temperature deSalas:2015glj ; Hasegawa:2019jsa sets constraints on the space of model parameters.
@@
    Fluctuations that are significantly amplified can potentially probe the dangerous region above the instability scale $\mu_{\rm inst}$ Tang:2013bz ; Espinosa:2015qea , defined by the Higgs self-coupling running to negative values, i.e. $\lambda(\mu_{\rm inst})=0$. Requiring the classical stability of our Universe imposes constraints on the top quark mass, as this is the major actor in the negative running of $\lambda$ Bezrukov:2014bra ; Hiller:2024zjp . [^3] The values of the top quark masses used in this work refer to the top quark pole mass, $m_{t}^{\text{pole}}$, which differs from the Monte Carlo reconstructed mass, $m_{t}^{\text{rec}}$, typically extracted using event generators like PYTHIA or HERWIG. The difference between these quantities is approximately 11112222 GeV. Recent measurements confirm that $m_{t}^{\text{pole}}$ is consistently lower than $m_{t}^{\text{rec}}$, with $m_{t}^{\text{pole}}=170.5\pm 0.8\,\text{GeV}$ CMS:2019esx and $m_{t}^{\text{rec}}=171.77\pm 0.37\,\text{GeV}$ CMS:2023ebf .
@@
    The generation of large spatial gradients leads inevitably to the production of a Stochastic Gravitational-Wave Background (SGWB). The signal is generated as the classicalised Higgs fluctuations enter the post-tachyonic non-linear dynamics, when the self-scattering process fragments the homogeneous patches inside the Hubble volume Bettoni:2024ixe .
[…]
@@
We are interested in exploring a large range of non-minimal coupling parameters $\xi$, inflationary scales $\mathcal{H}_{\rm inf}$ and top quark masses $m_{t}$ to identify the different scenarios that can be investigated by GW detection experiments. Our semi-analytical approach makes use of parametric formulas derived in previous works that condense numerous fully-fledged 3+1 classical lattice simulations into a few simple expressions Laverda:2023uqv ; Bettoni:2024ixe . We also perform three high-resolution benchmark simulations in the extended scenario involving non-renormalisable operators to check the implications for the GW signal.
The present work is laid out as follows. We define the setup of the model including the Higgs effective potential and the tachyonic dynamics in Section 2, followed by the general study of the GW spectrum in Section 3. This picture is applied to the SM-Higgs scenario in Section 4, focussing on the differences between the cases of absolute stability, instability, and independent heating sector. Finally, in Section 5 we explore the possibility of the effective potential being stabilised by new physics at high scales. Section 6 contains some final remarks and ideas for further development.
2 Effective Higgs potential with non-minimal curvature coupling

### v2

# Higgs-Induced Gravitational Waves: the Interplay of Non-Minimal Couplings, Kination and Top Quark Mass

@@ Abstract
We explore a minimal scenario where the sole Standard-Model Higgs is responsible for reheating the Universe after inflation, produces a significant background of gravitational waves and maintains the full classical stability of the electroweak vacuum. As the Higgs self-coupling runs toward negative values at high energy scales, a non-minimal interaction with curvature during a stiff background expansion era drives the Higgs fluctuations closer to the instability scale. This curvature-induced tachyonic instability leads to an intense production of Higgs particles, accompanied by a stochastic gravitational-wave background. The characteristic features of such signal can be directly correlated to the inflationary scale, the non-minimal coupling parameter and the top-quark Yukawa coupling. We distinguish between three possible scenarios: absolute stability with low top-quark masses, potential vacuum instability, and absolute stability with new physics above the instability scale. Our findings suggest that the detection of a peaked background of gravitational waves together with its inflationary tail has the potential to unveil the features of the Higgs effective potential at very high energy scales while providing a minimal explanation for the reheating phase and the emergence of the Standard-Model plasma in the early Universe. Unlike other studies in the literature, the generation of gravitational waves in our scenario does not depend on the quantum instability of the Standard Model vacuum.
@@ 1 Introduction
In recent years, the steady advancement of particle-collider experiments at the Large Hadron Collider has explored the Higgs sector up to energies of $13.6\times 10^{3}\;\rm GeV$. The several orders of magnitude separating this energy range from plausible inflationary scales up to $10^{16}\textrm{ GeV}$ leave us wondering about new physical phenomena in the Higgs sector that might have taken place in the early Universe. What role does the Higgs play in the history of the early Universe? One possibility is that it dominates the energy budget at very high energy scales. This assumption identifies the Higgs field with the scalar degree of freedom responsible for inflation, giving rise to the Higgs inflation paradigm Bezrukov:2007ep ; Bauer:2008zj ; Garcia-Bellido:2008ycs ; Barbon:2009ya ; Bezrukov:2009db ; Bauer:2010jg ; Burgess:2010zq ; Bezrukov:2010jz ; Giudice:2010ka ; Bezrukov:2014bra ; Hamada:2014wna ; George:2015nza ; Repond:2016sol ; Fumagalli:2016lls ; Bezrukov:2017dyv ; Rubio:2019ypq ; Shaposhnikov:2020fdv ; Dux:2022kuk ; Poisson:2023tja (for a review see Rubio:2018ogq ). In the present work, we will consider the opposite alternative, namely that of a spectator Higgs in the early Universe Markkanen:2017dlc ; Markkanen:2017edu ; Herranen:2014cua ; Herranen:2015ima ; Figueroa:2015rqa ; Opferkuch:2019zbd ; Laverda:2024qjt . In this scenario, the Higgs field remains an energetically-subdominant quantum field during and shortly after inflation, with a negligible backreaction on the gravitational background. In general, the whole Standard Model (SM) sector coupled to the Higgs field can be considered to be in its vacuum state, while the Universe is still far from achieving a thermalised state.
@@
In this context, non-minimal gravitational interactions can induce substantial new-physics effects within the Higgs sector. Studies of quantum fields in curved spacetime reveal in fact that tree-level operators associated with these interactions are naturally generated via the renormalisation of the energy-momentum tensor Birrell:1982ix . Assuming the corresponding coupling constants to vanish is therefore an ad hoc prescription valid only at a specific energy scale, since such a choice is not preserved under radiative corrections as first noted in Chernikov:1968zm ; Callan:1970ze ; Tagirov:1972vv . With these considerations, the physics of the Higgs field in the early Universe is dictated by an effective Lagrangian density
@@
$\mathcal{L}=\left(\frac{M^{2}_{P}}{2}-\xi H^{\dagger}H\right)R-g^{\mu\nu}(D_{\mu}H)^{\dagger}(D_{\nu}H)-\lambda\left(H^{\dagger}H-\frac{v^{2}_{{}_{\hskip-0.85355pt\textrm{EW}}}}{2}\right)^{2}+\mathcal{L}_{\rm SM}+\mathcal{L}_{\rm BSM}\,$, (1)
with the potential for the Higgs doublet $H$ depending on the quartic self-coupling $\lambda$ and the electroweak vacuum expectation value $v_{{}_{\hskip-0.85355pt\textrm{EW}}}$. The additional term $\mathcal{L}_{\rm SM}$ encompasses all SM interactions beside the Higgs-field kinetic and potential terms, while $\mathcal{L}_{\rm BSM}$ accounts for contributions from Beyond the Standard-Model (BSM) physics at high energy scales. In particular, $\mathcal{L}_{\rm BSM}$ includes the kinetic and potential terms of an inflationary degree of freedom $\phi$, which is assumed to dominate the energy content of the Universe during its earliest stages of evolution. In this work, we will remain agnostic about the specific inflationary framework governing the spectator Higgs field dynamics. Our only requirement is that inflation is followed by a kination or stiff expansion era, with the end of slow-roll resulting in a kinetically-dominated Universe with effective equation-of-state parameter $w_{\phi}=p_{\phi}/\rho_{\phi}\simeq 1$}. Such a sequence of background expansion phases can be found typically in models of quintessential inflation, where a steep potential interpolates between two plateaus at high and low scales in order to describe the inflationary and cosmological-constant epochs with a single scalar degree of freedom Peebles:1998qn ; Spokoiny:1993kt ; Bettoni:2021qfs . The subdominant Higgs field is assumed to be coupled only indirectly through gravitational interactions to the inflationary sector, with suppressed tree-level interactions in the Lagrangian (1). These features can be readily achieved by assuming a shift-symmetric inflaton field, such as in the case of variable gravity settings Wetterich:1987fm ; Wetterich:1994bg ; Rubio:2017gty ; Wetterich:2019qzx , which gives rise to an almost massless inflaton in the asymptotic regimes. However, other scenarios can include a inflation-to-kination transition as well, including asymmetric $\alpha$-attractors models Akrami:2017cir ; Dimopoulos:2017zvq ; Dimopoulos:2017tud ; Garcia-Garcia:2018hlc , and axion-like models Gouttenoire:2021jhk .
The transition from inflation to kination triggers a so-called Hubble-Induced Phase Transition (HIPT) Laverda:2023uqv ; Laverda:2024qjt ; Bettoni:2019dcw ; Bettoni:2021zhq ; Bettoni:2018utf ; Bettoni:2018pbl ; Mantziris:2024uzz ; Kierkla:2023uzo . More precisely, in a Friedmann–Lemaître–Robertson–Walker (FLRW) spacetime with metric tensor $g_{\mu\nu}={\rm diag}(-1,\,a^{2}(t)\delta_{ij})$, the gravitational contribution to the Higgs mass is determined by the Ricci or curvature scalar
$R=3(1-3w)\mathcal{H}^{2}\;,$ (2)
with $\mathcal{H}=\dot{a}/a$ the Hubble function and $w$ the global equation-of-state parameter, which, within the spectator field approximation, can be safely identified with that of the inflation field, $w\simeq w_{\phi}$. A variation of the background expansion rate of the Universe impacts therefore the Higgs effective mass, making it possible for new phases to appear in its effective potential. In particular, for inflationary scenarios involving a stiff expansion era following the end of inflation, the curvature scalar experiences a transition from positive (w=1w=-1) to negative values ($w=1$).
@@
During inflation, the large Hubble-induced mass suppresses the Higgs quantum fluctuations, both at super-horizon and sub-horizon scales, while preventing the generation of large isocurvature fluctuations Laverda:2024qjt ; Herranen:2014cua ; Herranen:2015ima ; Kohri:2016wof . At the onset of kination, however, the large negative gravitational mass term induces a spontaneous symmetry-breaking with Hubble-dependent true vacua. The Higgs quantum modes then experience an exponential tachyonic amplification that quickly builds up the occupation number of infrared (IR) modes, followed by a cascade effect that turbulently populates ultraviolet (UV) modes as well.  [^1] It is worth noting that this dynamic is significantly different from the well-known scenario of a resonantly-excited Higgs field interacting with an oscillating background Figueroa:2015rqa ; Figueroa:2016dsc ; Figueroa:2017slm ; Figueroa:2016wxr ; Herranen:2015ima ; Mantziris:2020rzh ; Mantziris:2021oah ; Mantziris:2021zox ; Mantziris:2022bfe ; Mantziris:2022fuu ; Mantziris:2023xsp ; Kohri:2016wof ; Postma:2017hbk ; Ema:2017loe ; Enqvist:2016mqj ; Ema:2016kpf . It is precisely these amplified fluctuations that provide us with a way to study the post-inflationary spectator Higgs as they lead to three main outcomes:
@@
    The interplay between kination and non-perturbative tachyonic particle production allows the spectator Higgs field itself to (re)heat the Universe, even without direct couplings of the inflaton field to matter or a full depletion of the inflaton condensate. Requiring the Universe to be radiation-dominated at temperatures higher than the Big-Bang-Nucleosynthesis (BBN) temperature deSalas:2015glj ; Hasegawa:2019jsa sets constraints on the space of model parameters.
@@
    Fluctuations that are significantly amplified can potentially probe the dangerous region above the instability scale $\mu_{\rm inst}$ Tang:2013bz ; Espinosa:2015qea , defined by the Higgs self-coupling running to negative values, i.e. $\lambda(\mu_{\rm inst})=0$. Requiring the classical stability of our Universe imposes constraints on the top-quark mass, as this is the major actor in the negative running of $\lambda$ Bezrukov:2014bra ; Hiller:2024zjp .
@@
    The generation of large spatial gradients leads inevitably to the production of a Stochastic Gravitational-Wave Background (SGWB). The signal is generated as the classicalised Higgs fluctuations enter the post-tachyonic non-linear dynamics, when the self-scattering process fragments the homogeneous patches generated by the tachyonic instability inside the Hubble volume Bettoni:2024ixe .
Figure 1: Higgs effective potential with a non-minimal gravitational coupling at the onset of kination for three prototypical scenarios: absolute stability (green line, $m_{t}=170.8\textrm{ GeV}$), instability (red line, $m_{t}=171.3\textrm{ GeV}$), stability with new physics as sextic operators at a typical scale $\Lambda=10^{15}\textrm{ GeV}$} (blue line, $m_{t}=171.3\textrm{ GeV}$). The cut-off of the effective gravitational theory is given by $\Lambda_{\xi}=9.2\times 10^{15}\textrm{ GeV}$. The cosmological parameters have been set to $\nu=20$ and $\mathcal{H}_{\rm kin}=10^{9}\textrm{ GeV}$.
The SGWB signal generated by HIPTs offers a unique window into the post-inflationary physics of the spectator Higgs field. In particular, the specific features of the Gravitational-Wave (GW) spectrum allow us to explore the running of the Higgs self-coupling at high energy scales, by indirectly measuring the top-quark mass. Three possible scenarios can occur. For top-quark masses below the critical value of $170.886\textrm{ GeV}$ that guarantees absolute stability, the GW signal can peak at $h^{2}\Omega_{\rm GW,0}\sim 10^{-9}$ and its typical frequency $f_{\rm GW,0}\sim 10^{10}\textrm{ Hz}$} is tightly linked to the top-quark mass. For higher top-quark masses, the constraint on the stability of the electroweak vacuum bounds the inflationary scale from above, thus limiting the amplitude of the signal to $h^{2}\Omega_{\rm GW,0}\lesssim 10^{-9}$. Notably, this GW background does not rely on the Higgs field crossing the instability scale, as considered, for instance, in Espinosa:2018eve . In the presence of new physics around the instability scale, a prototypical sextic operator in the Higgs potential can restore its absolute stability, see Figure 1. In this case, a high-scale phase transition is still compatible with stability even for high inflationary scales or values of the top-quark mass above the critical one.
@@
Exploring a large range of non-minimal coupling parameters $\xi$, inflationary scales $\mathcal{H}_{\rm inf}$, and top-quark masses $m_{t}$, allows us to identify telltale features in the GW spectrum. Our semi-analytical approach makes use of parametric formulas derived in previous works that condense numerous fully-fledged 3+1 classical lattice simulations into a few simple expressions Laverda:2023uqv ; Bettoni:2024ixe . We also perform three high-resolution benchmark simulations in the scenario involving non-renormalisable operators to check the implications for the GW signal.
This work is structured as follows. Section 2 provides an overview of the Higgs field’s behaviour during a HIPT transition, with particular focus on the validity of the spectator field approximation and the consistency of the effective field theory approach. In Section 3, we examine the impact of radiative corrections, emphasising the connection between the key cosmological parameters governing (re)heating dynamics and low-energy Standard Model inputs such as the top-quark mass. Section 4 then presents a general analysis of the gravitational wave spectrum produced by HIPTs, which is further applied to the Standard Model Higgs scenario in Section 5. This discussion highlights the differences between cases of absolute stability, instability, and an enhanced (or independent) heating sector. In Section 6, we explore the stabilisation of the effective potential by new physics at high energy scales. Finally, Section 7 offers concluding remarks and outlines directions for future research.
2 The spectator Higgs field approximation
To qualitatively discuss the general behaviour of the Higgs field during a HIPT Bettoni:2019dcw ; Laverda:2023uqv ; Bettoni:2021zhq and assess the validity of the spectator field approximation, it suffices to focus on the portion of (1) that involves the inflaton field $\phi$ and the subdominant Higgs field in the unitary gauge $H=(0,\,h/\sqrt{2})^{T}$, namely
${\cal L}=\frac{M^{2}_{P}}{2}R+\mathcal{L}_{\phi}-\frac{1}{2}{\partial^{\mu}}{h}\partial_{\mu}h-\frac{1}{2}\xi Rh^{2}-\frac{\lambda}{4}h^{4}\,$, (3)
where we have intentionally omitted the Higgs vacuum expectation value $v_{{}_{\hskip-0.85355pt\textrm{EW}}}$, as this will not play any essential role at the high energy scales considered in this paper. The dynamics of the Higgs field is governed by the Klein–Gordon equation,
$\Box\,h-\xi Rh-\lambda h^{3}=0\,$, (4)
along with the energy density and pressure following from the $00$ and $ii$ components of the associated energy-momentum tensor,
$T^{(h)}_{\mu\nu}=\partial_{\mu}h\partial_{\nu}h-g_{\mu\nu}\left[\frac{1}{2}\partial^{\rho}h\partial_{\rho}h+\frac{\lambda}{4}\,h^{4}\right]+\xi\Big[G_{\mu\nu}+\left(g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}\right)\Big]h^{2},$ (5)
with $\nabla_{\mu}$ the covariant derivative, $G_{\mu\nu}$ the Einstein tensor, and $\Box=\nabla^{\lambda}\nabla_{\lambda}$ defines the d’Alambertian operator. Notably, the contraction of $T^{(h)}_{\mu\nu}$ with arbitrary timelike vectors $u^{\mu}$ and $u^{\nu}$ is not necessarily positive definite at all times Ford:1987de ; Ford:2000xg ; Bekenstein:1975ww ; Flanagan:1996gw , leading to a violation of the weak energy condition $T^{(h)}_{\mu\nu}u^{\mu}u^{\nu}\geq 0$ within this particular sector.